# FROST-CLUSTERS – I. Hierarchical star cluster assembly boosts intermediate-mass black hole formation

### v1

@@ Acknowledgments
The authors thank Francesco Paolo Rizzuto and Jacob Kosowski from the Theoretical Extragalactic Group of the University of Helsinki, Finland. The numerical simulations were performed using facilities hosted by the Max Planck Computing and Data Facility (MPCDF) and the Leibniz Supercomputing Centre (LRZ), Germany. TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS".
@@ B.3.2 BH-BH mergers
A BH-BH occurs in the current version of BIFROST based on the radius of the innermost stable circular orbit of a zero-spin BH defined as $r_{\mathrm{ISCO}}=6Gm_{\bullet}/c^{2}$. The BH pair collision criterion is thus $r_{\mathrm{sep}}<\max{(r_{\mathrm{ISCO,p}},r_{\mathrm{ISCO,s}})}$. The mass, spin and recoil velocity of the remnant BH is calculated using the fitting formulas to numerical relativity simulations by Zlochower & Lousto (2015). Examples for recoil kick velocities for various BH binary mass ratios and spins are elaborated in section 2.4. The gravitational-wave mass loss is typically of the order of $\sim 5\%$ of the original total mass of the system (e.g. Jiménez-Forteza et al. 2017).

### v2

@@ Acknowledgments
The authors thank the reviewer for a constructive referee report. We also thank Francesco Paolo Rizzuto and Jacob Kosowski from the Theoretical Extragalactic Group of the University of Helsinki, Finland. The numerical simulations were performed using facilities hosted by the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany. TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS".
@@ B.3.2 BH-BH mergers
A BH-BH occurs in the current version of BIFROST based on the radius of the innermost stable circular orbit of a zero-spin BH defined as $r_{\mathrm{ISCO}}=6Gm_{\bullet}/c^{2}$. The BH pair collision criterion is thus $r_{\mathrm{sep}}<\max{(r_{\mathrm{ISCO,p}},r_{\mathrm{ISCO,s}})}$. The mass, spin and recoil velocity of the remnant BH is calculated using the fitting formulas to numerical relativity simulations by Zlochower & Lousto (2015). The natal BH spins are calculated using the Geneva model (Belczynski et al., 2020) as in their Eq. (3). Examples for recoil kick velocities for various BH binary mass ratios and spins are elaborated in section 2.4. The gravitational-wave mass loss is typically of the order of $\sim 5\%$ of the original total mass of the system (e.g. Jiménez-Forteza et al. 2017).